OpenStudy (anonymous):

which equation describes a parabola -2y^2+10x+9y=-11 -4x^2+10x-4y^2-2y+9=0 3y^2+10y-8x^2-4x-2=0 -2x^2+10x-4y^2+9y=2 help?!

OpenStudy (anonymous):

-2y^2+10x+9y=-11 (parabola) -4x^2+10x-4y^2-2y+9=0 (circle) 3y^2+10y-8x^2-4x-2=0 (hyperbola) -2x^2+10x-4y^2+9y=2 (ellipse)

OpenStudy (anonymous):

how is a a parabola ? @ilfy214

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=-2y%5E2%2B10x%2B9y%3D-11 look at the gaph

OpenStudy (anonymous):

These are the correct equations, right?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

Your welcome! :)